A metre rule is supported on a knife edge placed at the 40cm Mark.it is found that the metro rule balance when a mass of 45g is suspended from the 15cm .if the centre of mass of the the metre rule is at the 52.5cm mark.calculate the mass of metre rule
1
Expert's answer
2020-11-25T10:50:39-0500
"m(40-15)=M(52.5-45)\\45(40-15)=M(52.5-40)\\M=90\ g"